Reports 1-1 of 1 Clear search Modify search
mark.barton - 18:59 Tuesday 22 November 2016 (2182) Print this report
BS Test Hang

With Enzo.


  • We did a free swing test to see if everything was OK after the earthquake (at 5:59 am Japan time this morning) and to check that there weren't modes below the frequency used for the fixed-sine tests. Everything looked OK and the lowest frequency was a yaw mode near 0.4 Hz. See attached.
  • Data and plot are at /kagra/Dropbox/Subsystems/VIS/TypeBData/BSTestHang/IRM/2016-11-22-BSTH-IRM-DTT-dangle-01.*. 
  • We fixed more stuff on the VIS_CUST_TYPEB_OVERVIEW.adl and related screens:
    • The indicator for the front-end computer state word had been hard-wired to FEC 38, so we made it generic. (The FEC number for the PR2 front end currently being used is 96, but this will change when we get the actual BS rack, and the code should work for SRx as well.)
    • The links to the front-end computer GDS_TP screens were broken because "sus"/"SUS" had not been changed to "vis"/"VIS" in several places, and because the variables "comp"/"COMP" were set to "vis"/'VIS" not "pr2"/"PR2" in visbs_overview_macro.txt .
    • The VIS_CUST_TYPEB_ADC_MONITOR.ADL screen had bad channels with LIGO-style names.
    • The VIS_CUST_TYPEB_DAC_MONITOR.adl screen only supported the TM and had various non-generic items.
    • The variables for channel numbers in visbs_overview_macro.txt were too cryptic, so we changed them TH1=0_0 -> TMH1=0_0 etc in the macro file and screens that used them. (The TM ones will change again to OP when we properly implement Okutomi-san's naming scheme.)
  • We repeated the fixed-sine tests from yesterday (again with a frequency of 0.3 Hz) to check that the actuator polarities are now properly corrected for. Data is /kagra/Dropbox/Subsystems/VIS/TypeBData/BSTestHang/IRM/2016-11-22-BSTH-IRM-DTT-IM-??-01.xml.
  • With gains of +1, -1, -1, -1, -1, -1, phase angles are now:
    • V1: 1.5°
    • V2: 2.7°
    • V3: 0.4°
    • H1: 7.1°
    • H2: 1.7°
    • H3: 2.1°
  • We did a new snapshot with the new gains in.
Non-image files attached to this report
Comments to this report:
ayaka.shoda - 12:40 Wednesday 23 November 2016 (2184) Print this report

Hmm... Actually, they seem weird to me. Please correct me if I am wrong.

* IM YAW resonant frequency
In Type-Bp, IM resonant frequency in YAW is 0.14 Hz. (See k-log 1798)
While the PR maraging wire diameter (D_p) is 2.5, BS maraging wire diameter (D_b) is 3.6.
And the PR maraging wire length (L_p) is 598.5, BS maraging wire length (L_b) is 625.
The spring constant in the torsional mode is relative to D^4/L.
Ignoring the difference of the moment of inertia, the resonant frequency of the BS IM YAW mode is 0.14*(D_b/D_p)^2*(L_b/L_p)^(1/2) ~ 0.28 Hz.
More than that, considering the moment of inertia of the BS payload is larger than that of the PR payload, the YAW mode resonant frequency may be lower than 0.28 Hz.
Therefore, I think the resonance of 0.4 Hz for IM YAW is too high.

* Decoupling
Also, I think the signal decoupling is quite bad.
I think that the amount of displacement that each OSEM is monitoring is different too much.
As you can see the k-log 1798, the background OSEM signal levels are alsmot same in each signals.
(Note that 1798 is observed in the vacuum tank. The displacement itself can be larger than that.)

So... I recommend you to check if nothing is touching again, and compare the resonant frequencies with the calculation.

mark.barton - 11:01 Friday 25 November 2016 (2190) Print this report

To Shoda-san:

Thank you for pointing this out - I had missed it. You're right: the different magnitudes are very suspicious. Because the H1/H2/H3/V2 are suppressed, but V1 and V3 are relatively free, the primary suspect is probably some vertical stop screw near V2. It looks like I may have to go to Kamioka to take a look.

Search Help